OpenStudy (anonymous):

Will medal! Using a directrix of y = 2 and a focus of (3, -4), what quadratic function is created?

OpenStudy (anonymous):

Let me show you how (y+2)2=(x−1)2+(y−6)2 after some simplification you get−x2+2x−33=−16yy=116(x2−2x+33)=116(x2−2x+1+32)=116((x−1)2+32)=116(x−1)2+2